Measurement of elastic ω photoproduction at HERA

H1 Collaboration

DESY, Hamburg, Germany

The reaction γp → ωp (ω → π^{0}γ and π^{0} → γγ) is studied at photon-proton centre-of-mass energies around 200 GeV using the H1 detector at HERA. Measurements are presented of the photon-proton cross section for elastic ω production and of the differential cross section dσ/d|t|. The t-dependence is found to follow an exponential distribution of the form dσ/d|t| ∝ exp −b|t|. The slope parameter b is determined and is compared with the measurements of other elastic meson production processes.